DIFFERENTIAL GEOMETRY/PDE SEMINAR

Thursday, April 22, 2004 Padelford C-36 1:30 pm

Ricci flow preserves the isometry group

Peng Lu (University of Oregon)

By the uniqueness of the Ricci flow on closed manifolds that any isometry of the initial Riemannian manifold is preserved by the Ricci flow. A. Fischer raised: is the whole isometry group preserved by the Ricci flow on closed manifolds? The answer is affirmative, in this talk we will present a proof.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.